OpenStudy (anonymous):

The sum of two numbers is 112. If three times the smaller number is sibtracted from the larger number, the result is 12. Find the two numbers.

OpenStudy (anonymous):

25 smaller 87 larger?

myininaya (myininaya):

\[x+y=112\] from first sentence ----------------- Assume x is the smallest! \[y-3x=12\] from second sentence

myininaya (myininaya):

\[y=3x+12 \text{ from second equation}\] now plug this into the first equation and solve for x

myininaya (myininaya):

\[x+(3x+12)=112\] \[4x+12=112\] \[4x=100\] x=25 then y=3(25)+12=75+12=87 so yes

OpenStudy (nenadmatematika):

did he ask three times smaller number ?

OpenStudy (nenadmatematika):

I got it...sorry :D